what are the steps to solve the algebraic expression: 3c+4h=115?
Answer by checkley77(12844)

3c+4h=115
Solving for c:
3c=115-4h
c=(115-11h)/3 ans.
Soling for h:
4h=115-3c
h=(115-3c)/4 ans.